# Higher-Rank Irreducible Cartesian Tensors for Equivariant Message Passing

### v1

@@ D.1 Description of the data sets
3BPA data set. The 3BPA data set comprises structures, energies, and atomic forces of a flexible drug-like organic molecule obtained from AIMD at various temperatures [48]. The training data set consists of 500 configurations sampled at 300 K, while three separate test data sets are obtained from AIMD simulations at 300 K, 600 K, and 1200 K. An additional test data set provides energy values along dihedral rotations of the molecule. This test directly assesses the smoothness and accuracy of the potential energy surface, influencing properties such as binding free energies to protein targets. Our models are trained using 450 and 50 configurations randomly sampled from the training data set using five random seeds, with further 50 configurations reserved for early stopping. Table 2 reports the root-mean-square errors (RMSEs) in total energies and atomic forces averaged over five independent runs, including the standard deviation between them.
@@
Acetylacetone data set. The acetylacetone data set includes structures, energies, and atomic forces of a small reactive molecule obtained from AIMD at various temperatures [32]. The training data set comprises configurations sampled at 300 K, while the test data sets are sampled at 300 K and 600 K. The generalization ability of final models is evaluated using an elevated temperature of 600 K and along two internal coordinates of the molecule: The hydrogen transfer path and a partially conjugated double bond rotation featuring a high rotation barrier. Our models are trained using 450 and 50 configurations randomly sampled from the training dataset using five random seeds, with further 50 configurations reserved for early stopping. Table 3 reports the root-mean-square errors (RMSEs) in total energies and atomic forces averaged over five independent runs, including the standard deviation between them.
D.2 Training details
@@
All ICTP and MACE models employed in this work are trained on a single NVIDIA A100 GPU with 80 GB of RAM. Training times for ICTP and MACE models are typically between 30 minutes and 15 hours, depending on the data set, the data set size, and the employed precision. We use double precision for 3BPA and AcAc data sets and single precision for rMD17, in line with the original experiments [33]. For all experiments, we use two message-passing layers and irreducible Cartesian tensors or spherical tensors of a maximal rank of $l_{\mathrm{max}}=3$ to embed the directional information of atomic distance vectors. For ICTP models with the full and symmetric product basis (ICTPfull and ICTPsym, respectively) and MACE, we employ 256 uncoupled feature channels. For ICTP models with the symmetric product basis evaluated in the latent feature space (ICTPsym+lt), we use 64 coupled feature channels for the Cartesian product basis and 256 for two-body features. Radial features are derived from eight Bessel basis functions with polynomial envelope for the cutoff with $p=5$ [70]. These features are fed into a fully connected NN of size $[64,64,64]$. We apply SiLU non-linearities to the outputs of the hidden layers [83, 84]. The readout function of the first message-passing layer is implemented as a linear layer. The readout function of the second layer is a single-layer fully connected NN with 16 hidden neurons. A cutoff radius of 5 Å is used across all data sets.
@@
All parameters of ICTP and MACE models are optimized by minimizing the combined squared loss on training data $\mathcal{D}_{\mathrm{train}}=\left(\mathcal{X}_{\mathrm{train}},\mathcal{Y}_{\mathrm{train}}\right)$, where $\mathcal{X}_{\mathrm{train}}=\{\mathcal{S}^{(k)}\}_{k=1}^{N_{\mathrm{train}}}$ and $\mathcal{Y}_{\mathrm{train}}=\{E_{k}^{\mathrm{ref}},\{\mathbf{F}_{u,k}^{\mathrm{ref}}\}_{u=1}^{N_{\mathrm{at}}}\}_{k=1}^{N_{\mathrm{train}}}$:
@@
$\mathcal{L}\left(\boldsymbol{\theta},\mathcal{D}_{\mathrm{train}}\right)=\sum_{k=1}^{N_{\mathrm{train}}}\Bigg{[}C_{\mathrm{e}}\Big{\lVert}E_{k}^{\mathrm{ref}}-E(\mathcal{S}^{(k)},\boldsymbol{\theta})\Big{\rVert}_{2}^{2}+C_{\mathrm{f}}\sum_{u=1}^{N_{\mathrm{at}}^{(k)}}\Big{\lVert}\mathbf{F}_{u,k}^{\mathrm{ref}}-\mathbf{F}_{u}\left(\mathcal{S}^{(k)},\boldsymbol{\theta}\right)\Big{\rVert}_{2}^{2}\Bigg{]}.$ (A5)
@@
When training ICTP models, we set $C_{\mathrm{e}}=1/N_{\mathrm{at}}^{(k)}$ and $C_{\mathrm{f}}=10~{}\text{\AA}^{2}$ to balance the relative contributions of total energies and atomic forces, respectively. For MACE, $C_{\mathrm{e}}=1/(B\times N_{\mathrm{at}}^{(k)})$ and $C_{\mathrm{f}}=1000/(B\times 3\times N_{\mathrm{at}}^{(k)})~{}\text{\AA}^{2}$ are used with $B$ denoting the batch size. Here, $E(\mathcal{S}^{(k)},\boldsymbol{\theta})$ and $\mathbf{F}_{u}\left(\mathcal{S}^{(k)},\boldsymbol{\theta}\right)$ are total energies and atomic forces predicted by ICTP or MACE.
@@
All models are trained for 2000 epochs using the AMSGrad variant of Adam [85], with default parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\varepsilon=10^{-8}$. We use a learning rate of 0.01 and a batch size of 5 for all experiments. For evaluations on the validation and test data sets, we use a batch size of 10. The learning rate is reduced using an on-plateau scheduler based on the validation loss with a patience of 50 and a decay factor of 0.8. We utilize an exponential moving average with a weight of 0.99 for evaluation on the validation set and for the final model. Additionally, in line with MACE [33], we apply exponential weight decay of $5\times 10^{-7}$ on the weights of Eqs. (6), (A1), and (A3). Furthermore, we incorporate a per-atom shift of total energies via the average per-atom energy over all the training configurations, including the energies of individual atoms for 3BPA and acetylacetone datasets. If no atomic energies are provided, as for rMD17, the per-atom shift is obtained by solving a linear regression problem [56]. Additionally, a per-atom scale is determined as the root-mean-square of the components of the forces over the training configurations.
D.3 Additional results
Extrapolation to out-of-domain data. Table A1 demonstrates total energy and atomic force RMSEs obtained for ICTP and MACE models trained using 50 configurations randomly drawn from the original data set. ICTP and MACE models perform similarly, considering the standard deviation obtained across five independent runs. However, ICTP models often have lower mean RMSE values compared to MACE.
@@ D.3 Additional results
Table A1: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 50 molecules randomly drawn from the training data set collected at 300 K, while additional 50 are used for early stopping. All values for ICTP and MACE models are obtained by averaging over five independent runs, with the standard deviation provided if available. Best performances, considering the standard deviation, are highlighted in bold.
@@
Figure A1: Potential energy profiles for three cuts through the 3BPA molecule’s potential energy surface. All models are trained using 450 configurations, and the remaining 50 are used for early stopping. The 3BPA molecule, including the three dihedral angles ($\alpha$, $\beta$, and $\gamma$), provided in degrees , is shown as an inset. The color code of the inset molecule is C grey, O red, N blue, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. Shaded areas denote standard deviations across five independent runs.
Figure A2 further investigates the generalization capabilities of ICTP models trained using 450 configurations, demonstrating potential energy profiles for the rotation around the C-C bond, i.e., the C-C-C-O dihedral angle ($\alpha$), and for the hydrogen transfer. The training data set encompasses dihedral angles less than $30^{\circ}$. Furthermore, the energy barrier of 1 eV is outside the energy range of the training data set obtained at 300 K. As for the hydrogen transfer, the training data does not contain transition geometries, but the reaction still occurs in a region that is not too far from the training data. Overall, ICTP models perform on par with MACE for predicting potential energy profiles for the rotation around the corresponding C-C bond and for the hydrogen transfer.
@@
Table A2: Energy (E) and force (F) root-mean-square errors (RMSEs) for the acetylacetone data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 50 configurations randomly drawn from the training data set collected at 300 K, while additional 50 are used for early stopping. All values for ICTP and MACE models are obtained by averaging over five independent runs. Best performances, considering the standard deviation, are highlighted in bold.
@@
Figure A2: Potential energy profiles of (a) the dihedral angle describing the rotation around the C-C bond and (b) hydrogen transfer between two oxygen atoms. All models are trained using 450 molecules, and the remaining 50 are used for early stopping. The acetylacetone molecule, including the dihedral angle in degrees describing the rotation around the C-C bond ($\alpha$), is shown as an inset in (a). The color code of the inset molecule is C grey, O red, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. The histograms demonstrate the distribution of dihedral angles and O-H distances in the training data. Shaded areas denote standard deviations across five independent runs.

### v2

@@ E.1 Description of the data sets
MD22 data set. The MD22 data set includes structures, energies, and atomic forces of seven molecular systems derived from AIMD simulations at elevated temperatures, spanning four major classes of biomolecules and supramolecules [48]. These molecular systems range from a small peptide containing 42 atoms to a double-walled nanotube comprising 370 atoms. Characterized by complex intermolecular interactions, this data set was designed to challenge short-range models. The training set sizes used in this study are consistent with those in the original publication [48], selected to ensure that the sGDML model stays within a target accuracy of 1 kcal/mol ($\approx 43.37$ meV). We randomly selected an additional set of 500 structures for each molecule in the data set for early stopping, while the remaining configurations were reserved for testing the final models. The corresponding training and validation data sets were randomly selected using three random seeds. Table A2 reports MAEs in total energies per atom and atomic forces averaged over three independent runs, including the standard deviation between them.
@@
3BPA data set. The 3BPA data set comprises structures, energies, and atomic forces of a flexible drug-like organic molecule obtained from AIMD at various temperatures [49]. The training data set consists of 500 configurations sampled at 300 K, while three separate test data sets are obtained from AIMD simulations at 300 K, 600 K, and 1200 K. An additional test data set provides energy values along dihedral rotations of the molecule. This test directly assesses the smoothness and accuracy of the potential energy surface, influencing properties such as binding free energies to protein targets. Our models are trained using 450 and 50 configurations randomly sampled from the training data set using five random seeds, with further 50 configurations reserved for early stopping. Table 2 reports the root-mean-square errors (RMSEs) in total energies and atomic forces averaged over five independent runs for a training set size of 450 structures, including the standard deviation between them. Table A3 presents the corresponding results obtained with a training set size of 50 structures.
@@
Acetylacetone data set. The acetylacetone data set includes structures, energies, and atomic forces of a small reactive molecule obtained from AIMD at various temperatures [32]. The training data set comprises configurations sampled at 300 K, while the test data sets are sampled at 300 K and 600 K. The generalization ability of final models is evaluated using an elevated temperature of 600 K and along two internal coordinates of the molecule: The hydrogen transfer path and a partially conjugated double bond rotation featuring a high rotation barrier. Our models are trained using 450 and 50 configurations randomly sampled from the training dataset using five random seeds, with further 50 configurations reserved for early stopping. Table 3 reports RMSEs in total energies and atomic forces averaged over five independent runs for a training set size of 450 structures, including the standard deviation between them. Table A5 presents the corresponding results obtained with a training set size of 50 structures.
Ta–V–Cr–W data set. The Ta–V–Cr–W data set includes 0 K energies, atomic forces, and stresses for binaries, ternaries, and quaternary and near-melting temperature properties in four-component disordered high-entropy alloys [50]. In total, this benchmark data set comprises 6711 configurations, with energies, atomic forces, and stresses computed at the density functional theory (DFT) level. More precisely, there are 5680 0 K structures: 4491 binary, 595 ternary, and 594 quaternary structures, along with 1031 structures sampled from MD at 2500 K. Structure sizes range from 2 to 432 atoms in the periodic cell. All models are trained using 5373 configurations, with 4873 used for training and 500 for early stopping. The remaining configurations are reserved for testing the models’ performance. The performance is evaluated separately using 0 K binaries, ternaries, quaternaries, and near-melting temperature four-component disordered alloys. The original data set provides ten different training-test splits. In our experiments, each training set is further split into training and validation subsets using a random seed. Furthermore, we use additional binary structures strained along the $[100]$ direction as a part of the test data set. Note that the final models, in this case, were obtained by training using the whole data set of 6711 configurations (training + test), 500 of which were reserved as a validation data set. Table A6 reports RMSEs in total energies per atom and atomic forces averaged over ten independent runs, including the standard deviation between them.
E.2 Training details
@@
All ICTP and MACE models employed in this work were trained on a single NVIDIA A100 GPU with 80 GB of RAM. Training times for ICTP and MACE models typically ranged from 30 minutes to a few days, depending on the data set, the data set size, and the employed precision. We used double precision for 3BPA and acetylacetone data sets and single precision for rMD17, in line with the original experiments [33]. Double precision was also used for MD22, while single precision was employed in our Ta–V–Cr–W experiments. Unless stated otherwise, we used two message-passing layers and irreducible Cartesian tensors or spherical tensors of a maximal rank of $l_{\mathrm{max}}=3$ to embed the directional information of atomic distance vectors.
For ICTP models with the full (ICTPfull) and symmetric (ICTPsym) product basis and MACE, we employ 256 uncoupled feature channels. Exceptions include our experiments with the 3BPA data set, aimed at investigating scaling and computational cost, and the Ta–V–Cr–W experiments, where we used eight and 32 feature channels, respectively. For ICTP models with the symmetric product basis evaluated in the latent feature space (ICTPsym+lt), we use 64 coupled feature channels for the Cartesian product basis and 256 for two-body features. Radial features are derived from eight Bessel basis functions with polynomial envelope for the cutoff with $p=5$ [60]. These features are fed into a fully connected NN of size $[64,64,64]$. We apply SiLU non-linearities to the outputs of the hidden layers [87, 88]. The readout function of the first message-passing layer is implemented as a linear layer. The readout function of the second layer is a single-layer fully connected NN with 16 hidden neurons. A cutoff radius of 5.0 Å is used across all data sets except MD22, where we used a cutoff radius of 5.5 Å for the double-walled nanotube and 6.0 Å for the other molecules in the data set.
@@
All parameters of ICTP and MACE models were optimized by minimizing the combined squared loss on training data $\mathcal{D}_{\mathrm{train}}=\left(\mathcal{X}_{\mathrm{train}},\mathcal{Y}_{\mathrm{train}}\right)$, where $\mathcal{X}_{\mathrm{train}}=\{\mathcal{S}^{(k)}\}_{k=1}^{N_{\mathrm{train}}}$ and $\mathcal{Y}_{\mathrm{train}}=\{E_{k}^{\mathrm{ref}},\{\mathbf{F}_{u,k}^{\mathrm{ref}}\}_{u=1}^{N_{\mathrm{at}}},\boldsymbol{\sigma}_{k}^{\mathrm{ref}}\}_{k=1}^{N_{\mathrm{train}}}$
@@
$\begin{split}\mathcal{L}\left(\boldsymbol{\theta},\mathcal{D}_{\mathrm{train}}\right)=\sum_{k=1}^{N_{\mathrm{train}}}\Bigg{[}C_{\mathrm{e}}\Big{\lVert}E_{k}^{\mathrm{ref}}-E(S^{(k)},\boldsymbol{\theta})\Big{\rVert}_{2}^{2}&+C_{\mathrm{f}}\sum_{u=1}^{N_{\mathrm{at}}^{(k)}}\Big{\lVert}\mathbf{F}_{u,k}^{\mathrm{ref}}-\mathbf{F}_{u}\left(S^{(k)},\boldsymbol{\theta}\right)\Big{\rVert}_{2}^{2}\\ &+C_{\mathrm{s}}\Big{\lVert}V_{k}\boldsymbol{\sigma}_{k}^{\mathrm{ref}}-V_{k}\boldsymbol{\sigma}\left(S^{(k)},\boldsymbol{\theta}\right)\Big{\rVert}_{2}^{2}\Bigg{]}.\end{split}$ (A5)
Here, $\boldsymbol{\sigma}_{k}^{\mathrm{ref}}$ is the stress tensor defined as $\boldsymbol{\sigma}=\frac{1}{V}\left.\nabla_{\boldsymbol{\epsilon}}E\right|_{\boldsymbol{\epsilon}=\mathbf{0}}$, where $E$ denotes total energy after a strain deformation with symmetric tensor $\boldsymbol{\epsilon}\in\mathbb{R}^{3\times 3}$ and $V$ is the volume of the periodic box.
@@
When training ICTP models on rMD17, 3BPA, and acetylacetone data sets, we neglected the stress loss and set $C_{\mathrm{e}}=1/N_{\mathrm{at}}^{(k)}$ and $C_{\mathrm{f}}=10~{}\text{\AA}^{2}$ to balance the relative contributions of total energies and atomic forces, respectively. For MACE and 3BPA/acetylacetone, $C_{\mathrm{e}}=1/(B\times N_{\mathrm{at}}^{(k)})$ and $C_{\mathrm{f}}=1000/(B\times 3\times N_{\mathrm{at}}^{(k)})~{}\text{\AA}^{2}$ were used with $B$ denoting the batch size. For ICTP models trained on the MD22 data set, we set $C_{\mathrm{e}}=10/N_{\mathrm{at}}^{(k)}$ and $C_{\mathrm{f}}=1~{}\text{\AA}^{2}$, using energies and forces in eV and eV/Å, respectively. For the Ta–V–Cr–W dataset, the stress loss was incorporated into the combined loss in Eq. (A5), along with the energy and force losses. For ICTP, we used $C_{\mathrm{e}}=1/N_{\mathrm{at}}^{(k)}$, $C_{\mathrm{f}}=0.01~{}\text{\AA}^{2}$, and $C_{\mathrm{s}}=0.001/N_{\mathrm{at}}^{(k)}$ to balance the relative contributions of total energies, atomic forces, and stresses, respectively. For MACE, we chose $C_{\mathrm{e}}=1/(B\times N_{\mathrm{at}}^{(k)}\times N_{\mathrm{at}}^{(k)})$, $C_{\mathrm{f}}=1/(B\times 3\times N_{\mathrm{at}}^{(k)})~{}\text{\AA}^{2}$, and $C_{\mathrm{s}}=0.05/(B\times 9\times N_{\mathrm{at}}^{(k)}\times N_{\mathrm{at}}^{(k)})$. Here, $E(\mathcal{S}^{(k)},\boldsymbol{\theta})$, $\mathbf{F}_{u}\left(\mathcal{S}^{(k)},\boldsymbol{\theta}\right)$, and $\boldsymbol{\sigma}\left(S^{(k)},\boldsymbol{\theta}\right)$ are total energies, atomic forces, and stresses predicted by ICTP or MACE.
All models for rMD17, 3BPA, and acetylacetone were trained for 2000 epochs using the AMSGrad variant of Adam [89], with default parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\varepsilon=10^{-8}$. For MD22 and Ta–V–Cr–W, all models were trained for 1000 epochs. For rMD17, 3BPA, and acetylacetone data sets, we used a learning rate of 0.01 and a batch size of 5. For MD22 and Ta–V–Cr–W, we again chose a learning rate of 0.01 but a mini-batch of 2 and 32, respectively. For evaluations on the validation and test data sets, we used a batch size of 10 for rMD17, 3BPA, and acetylacetone, while mini-batches of 2 and 32 were used for MD22 and Ta–V–Cr–W, respectively.
@@
The learning rate was reduced using an on-plateau scheduler based on the validation loss with a patience of 50 and a decay factor of 0.8 for all data sets except for MD22, for which we used a patience of 10. We utilize an exponential moving average with a weight of 0.99 for evaluation on the validation set and for the final model. Additionally, in line with MACE [33], we apply exponential weight decay of $5\times 10^{-7}$ on the weights of Eqs. (6), (A1), and (A3). Furthermore, we incorporate a per-atom shift of total energies via the average per-atom energy over all the training configurations, including the energies of individual atoms for 3BPA and acetylacetone datasets. If no atomic energies are provided, as for rMD17, MD22, and Ta–V–Cr–W, the per-atom shift is obtained by solving a linear regression problem [67]. Additionally, a per-atom scale is determined as the root-mean-square of the components of the forces over the training configurations.
E.3 Additional results
Table A1: Inference times and memory consumption as a function of the tensor rank $L$ and the correlation order $\nu$ for the 3BPA data set. All values for ICTP and MACE models are obtained by averaging over five independent runs. The standard deviation is provided if it is available. Best performances are highlighted in bold. Inference time and memory consumption are measured for a batch size of 10. Inference time is reported per structure in ms; memory consumption is provided for the entire batch in GB.
$L=1$ $L=2$ $L=3$
[…]
Inference times
$\nu=1$ 0.76 $\pm$ 0.17 1.02 $\pm$ 0.03 0.87 $\pm$ 0.18 1.38 $\pm$ 0.04 0.98 $\pm$ 0.26 1.88 $\pm$ 0.03
$\nu=2$ 0.59 $\pm$ 0.20 1.12 $\pm$ 0.03 1.03 $\pm$ 0.21 1.52 $\pm$ 0.05 1.34 $\pm$ 0.08 2.0 $\pm$ 0.10
$\nu=3$ 0.79 $\pm$ 0.22 1.23 $\pm$ 0.03 1.15 $\pm$ 0.08 1.67 $\pm$ 0.03 1.85 $\pm$ 0.13 2.23 $\pm$ 0.03
$\nu=4$ 0.94 $\pm$ 0.17 1.41 $\pm$ 0.11 1.31 $\pm$ 0.21 1.83 $\pm$ 0.01 2.07 $\pm$ 0.20 2.53 $\pm$ 0.01
$\nu=5$ 1.02 $\pm$ 0.17 1.52 $\pm$ 0.08 1.72 $\pm$ 0.07 2.26 $\pm$ 0.03 3.61 $\pm$ 0.02 OOM
$\nu=6$ 1.0 $\pm$ 0.07 1.77 $\pm$ 0.05 1.83 $\pm$ 0.16 27.85 $\pm$ 0.01 16.76 $\pm$ 0.35 OOM
Memory consumption
$\nu=1$ 0.05 $\pm$ 0.00 0.04 $\pm$ 0.00 0.08 $\pm$ 0.00 0.06 $\pm$ 0.00 0.21 $\pm$ 0.00 0.13 $\pm$ 0.00
$\nu=2$ 0.05 $\pm$ 0.00 0.04 $\pm$ 0.00 0.08 $\pm$ 0.00 0.07 $\pm$ 0.00 0.28 $\pm$ 0.09 0.13 $\pm$ 0.00
$\nu=3$ 0.05 $\pm$ 0.00 0.04 $\pm$ 0.00 0.10 $\pm$ 0.00 0.08 $\pm$ 0.00 0.51 $\pm$ 0.03 0.23 $\pm$ 0.00
$\nu=4$ 0.05 $\pm$ 0.00 0.05 $\pm$ 0.00 0.18 $\pm$ 0.08 0.30 $\pm$ 0.00 1.07 $\pm$ 0.10 4.16 $\pm$ 0.00
$\nu=5$ 0.05 $\pm$ 0.00 0.07 $\pm$ 0.00 0.35 $\pm$ 0.07 3.18 $\pm$ 0.00 5.07 $\pm$ 0.02 OOM
$\nu=6$ 0.11 $\pm$ 0.09 0.22 $\pm$ 0.00 0.93 $\pm$ 0.00 50.49 $\pm$ 0.00 28.48 $\pm$ 0.03 OOM
Scaling and computational cost. Table A1 provides the numerical results complementing Fig. 2 in the main text. All results in Table A1 and Fig. 2 were obtained using irreducible Cartesian tensors with a maximal rank of $l_{\mathrm{max}}=L$ to represent local atomic environments. Here, $L$ denotes the tensor rank of employed equivariant messages. We facilitated the exploration of larger $\nu$ values by setting the number of uncoupled feature channels to eight. In the MACE model, intermediate spherical tensors with a rank of $l>l_{\mathrm{max}}$ are used to construct the product basis. However, the pre-computation of generalized Clebsch–Gordan coefficients for $\nu>4$, in some cases, would require more than 2 TB of RAM. Therefore, in our experiments, we fixed the maximum rank of intermediate tensors to $l=l_{\mathrm{max}}$. We also used the full product basis for ICTP to calculate the same number of $\nu$-fold tensor products, i.e., $\mathcal{K}=\mathrm{len}\left(\eta_{\nu}\right)$, as used in MACE with $l=l_{\mathrm{max}}$. Finally, we note that ICTP and MACE use different approaches to optimize their runtimes; however, the scaling with respect to the tensor rank and the correlation order is independent of these optimization methods.
@@ E.3 Additional results
Table A2: Energy (E) and force (F) mean absolute errors (MAEs) for the MD22 data set.a E- and F-MAE are given in meV/atom and meV/Å, respectively. Results are shown for models trained using training set sizes defined in the original publication [48]. All values for ICTP models are obtained by averaging over three independent runs. We also use an additional subset of 500 configurations drawn randomly from the original data set for early stopping. The standard deviation is provided if available. Best performances, considering the standard deviation, are highlighted in bold.
[…]
Ac–Ala3–NHMe E 0.068 $\pm$ 0.000 0.068 0.085 0.348 0.064 0.105 0.116 0.403
F 3.28 $\pm$ 0.02 3.91 3.49 10.58 3.8 4.63 8.15 34.26
DHA E 0.080 $\pm$ 0.001 0.068 0.138 0.293 0.102 0.089 0.093 0.997
F 2.62 $\pm$ 0.01 2.59 2.19 10.49 2.8 3.17 5.24 32.52
Stachyose E 0.053 $\pm$ 0.001 0.053 0.070 0.22 0.062 0.124 0.069 1.995
F 2.51 $\pm$ 0.03 3.33 2.75 18.86 3.8 4.21 8.33 29.49
AT–AT E 0.057 $\pm$ 0.002 0.056 0.095 0.129 0.079 0.103 0.081 0.52
F 3.02 $\pm$ 0.07 3.39 4.16 9.37 4.3 4.13 8.83 29.92
AT–AT–CG–CG E 0.045 $\pm$ 0.001 0.042 0.055 0.127 0.058 0.145 0.076 0.52
F 3.37 $\pm$ 0.05 4.61 5.43 14.40 5.0 5.55 14.13 30.36
Buckyball catcher E 0.123 $\pm$ 0.001 0.124 0.117 0.112 0.141 0.154 0.152 0.343
F 3.58 $\pm$ 0.07 4.45 4.83 10.28 3.7 3.85 14.39 29.49
Double-walled E 0.352 $\pm$ 0.003b 0.214 0.140 0.116 0.194 0.259 0.173 0.468
nanotube F 12.64 $\pm$ 0.23b 14.71 11.91 31.53 12.0 14.87 43.5 22.55
a ICTP, MACE, Equiformer, Allegro, and TorchMD-Net are short-range models (i.e., they use local or semi-local atomic representations), ViSNet-LSRM and SO3krates include long-range information, and sGDML is a global model.
b For consistency, these results are obtained for relative energy and force weights of $C_{\mathrm{e}}=10/N_{\mathrm{at}}^{(k)}$ and $C_{\mathrm{f}}=1~{}\text{\AA}^{2}$, based on our experiments with the Ac–Ala3–NHMe molecule. For comparison, we also trained ICTP for 800 epochs with $C_{\mathrm{e}}=100/N_{\mathrm{at}}^{(k)}$ and $C_{\mathrm{f}}=1~{}\text{\AA}^{2}$, achieving MAEs of 0.209 $\pm$ 0.007 meV/atom for energy and 13.98 $\pm$ 0.30 meV/Å for force.
Molecular dynamics trajectories. Table A2 presents the energy and force mean absolute errors (MAEs) for the ICTP models trained on the MD22 data set. This data set was designed to challenge short-range potential models—i.e., those based on local or semi-local atomic representations—and includes large molecular systems with complex intermolecular interactions. We compare errors obtained for ICTP models with those of state-of-the-art approaches that incorporate global, short- and long-range information. From Table A2, we see that ICTP achieves accuracy in predicted energies and atomic forces on par with or better than state-of-the-art methods. The ICTP model uses a cutoff of 6.0 Å (5.5 Å for the double-walled nanotube), resulting in a receptive field of 12.0 Å (11.0 Å), considering the two message-passing layers. This receptive field is, in most cases, smaller than the diameter of molecular systems in the data set. Thus, ICTP is, at most, a semi-local potential model—similar to MACE, which used a 5.0 Å cutoff and two message-passing layers.
We chose a larger cutoff for ICTP than the one used by MACE motivated by the results in the recent work [7]. In particular, the authors compared MACE to VisNet-LSRM—the best model reported to date for MD22, which employs mixed short- and long-range information in their message passing—and reported that MACE can achieve lower force errors. However, VisNet-LSRM typically had lower energy errors, attributed to the improvement from considering atomic interactions beyond 10.0 Å. Using a larger cutoff radius for ICTP compared to MACE, we expected results closer to those of VisNet-LSRM. Table A2 shows that using a receptive field of 12.0 Å is, in most cases, sufficient to achieve accuracy in predicted energies close to the one obtained by VisNet-LSRM.
Table A3: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set (results for $N_{\mathrm{train}}=50$). E- and F-RMSE are given in meV and meV/Å, respectively. Results are shown for models trained using 50 molecules randomly drawn from the training data set collected at 300 K, with further 50 used for early stopping. All ICTP and MACE results are obtained by averaging over five independent runs, with the standard deviation provided if available. Best performances, considering the standard deviation, are highlighted in bold.
@@
Table A4: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set (results for $\nu=1$). E- and F-RMSE are given in meV and meV/Å, respectively. Results are shown for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, with further 50 used for early stopping. For all models, $\nu=1$ is used. Thus, the employed models rely exclusively on two-body interactions. All ICTP and MACE results are obtained by averaging over five independent runs. Best performances, considering the standard deviation, are highlighted in bold. Inference time and memory consumption are measured for a batch size of 100. Inference time is reported per structure in ms, while memory consumption is provided for the entire batch in GB.
@@
300 K E 12.90 $\pm$ 1.06 12.90 $\pm$ 1.06 14.97 $\pm$ 1.64 13.50 $\pm$ 1.71
F 29.90 $\pm$ 0.25 29.90 $\pm$ 0.25 30.93 $\pm$ 0.47 30.18 $\pm$ 0.38
600 K E 29.97 $\pm$ 0.94 29.97 $\pm$ 0.94 31.64 $\pm$ 1.83 31.32 $\pm$ 2.16
F 62.80 $\pm$ 0.45 62.80 $\pm$ 0.45 64.54 $\pm$ 0.95 63.04 $\pm$ 0.73
1200 K E 81.03 $\pm$ 1.64 81.03 $\pm$ 1.64 79.26 $\pm$ 4.66 81.54 $\pm$ 2.02
F 146.96 $\pm$ 1.30 146.96 $\pm$ 1.30 151.45 $\pm$ 5.21 149.44 $\pm$ 1.94
Dihedral slices E 22.84 $\pm$ 2.96 22.84 $\pm$ 2.96 29.16 $\pm$ 7.96 28.08 $\pm$ 4.04
F 48.82 $\pm$ 5.25 48.82 $\pm$ 5.25 52.53 $\pm$ 5.27 49.62 $\pm$ 2.92
Inference time 2.63 $\pm$ 0.02 2.62 $\pm$ 0.02 2.53 $\pm$ 0.01 2.96 $\pm$ 0.06
Memory consumption 32.57 $\pm$ 0.00 32.57 $\pm$ 0.00 32.34 $\pm$ 0.09 23.32 $\pm$ 0.00
Figure A1: Potential energy profiles for three cuts through the 3BPA molecule’s potential energy surface (results for $N_{\mathrm{train}}=450$). All models are trained using 450 configurations, and the remaining 50 are used for early stopping. The 3BPA molecule, including the three dihedral angles ($\alpha$, $\beta$, and $\gamma$), provided in degrees , is shown as an inset. The color code of the inset molecule is C grey, O red, N blue, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. Shaded areas denote standard deviations across five independent runs.
[…]
Table A5: Energy (E) and force (F) root-mean-square errors (RMSEs) for the acetylacetone data set (results for $N_{\mathrm{train}}=50$). E- and F-RMSE are given in meV and meV/Å, respectively. Results are shown for models trained using 50 configurations randomly drawn from the training data set collected at 300 K, with further 50 used for early stopping. All ICTP and MACE results are obtained by averaging over five independent runs. Best performances, considering the standard deviation, are highlighted in bold.
@@
Figure A2: Potential energy profiles of (a) the dihedral angle describing the rotation around the C-C bond and (b) hydrogen transfer between two oxygen atoms (results for $N_{\mathrm{train}}=450$). All models are trained using 450 molecules, and the remaining 50 are used for early stopping. The acetylacetone molecule, including the dihedral angle in degrees describing the rotation around the C-C bond ($\alpha$), is shown as an inset in (a). The color code of the inset molecule is C grey, O red, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. The histograms demonstrate the distribution of dihedral angles and O-H distances in the training data. Shaded areas denote standard deviations across five independent runs.